OpenStudy (anonymous):

I need some help with algebra 2 please!!!

OpenStudy (anonymous):

@absurdism

OpenStudy (anonymous):

sure, post your equations

OpenStudy (anonymous):

x+y equals 3

OpenStudy (anonymous):

y-0 equals 1(x+3)

OpenStudy (anonymous):

y equals 2x+(-4)

OpenStudy (anonymous):

the DOMAIN set describes what numbers CAN go into the function without giving you an erroneous expression (a bad expression, like division by 0) the RANGE set describes what numbers are made as a RESULT of the domain inputs.

OpenStudy (anonymous):

and y equals -1x+3

OpenStudy (anonymous):

x + y = 3 rearrange this into slope-intercept form y = -x + 3 are there any numbers we can or cannot put into this? are there any numbers that we can't make with this function?

OpenStudy (anonymous):

i understand that but im not really sure what to do? like i have no idea what those definitons mean, like my mind cant grasp the concept

OpenStudy (anonymous):

@skatergurl916

OpenStudy (anonymous):

Um, i dont think we can put anything into it?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

like, by "put in" i mean substitute x for a number

OpenStudy (anonymous):

no

OpenStudy (anonymous):

maybe 1?

OpenStudy (anonymous):

i dont get this very well but theres a site that will give u all the answers u need

OpenStudy (anonymous):

how so? you can put any number into this function and it'll be fine, look y = -0 + 3 y = -1 + 3 y = -99999 + 3 it's all fine

OpenStudy (anonymous):

but doesnt that change the slope?

OpenStudy (jhannybean):

Graph your equation, maybe that will help.

OpenStudy (anonymous):

oh so does that mean i can put any number for the slope because of the domain?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

The domain is all the x values from negative infinity to positive infinity, so is there any place where the domain would not exist for this line?

OpenStudy (anonymous):

i dont understand the question on my assignment.. would any restriction apply to the domain and range in the equations i created

OpenStudy (jhannybean):

Similar for range, range is all the y values across the y-axis. Is there any place where the y values would be discontinuous? No, for both of these cases.

OpenStudy (anonymous):

ohh, so there are endless possiblies for this equation?

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

okay so i should write something like... when changed to slope intercept form .. the domain and range are endless?

OpenStudy (anonymous):

what about y-0 equals 1(x+3) would the possiblies be endless there as well.. sorry im a little confused.

OpenStudy (jhannybean):

the domain and range range from \((-\infty \ , +\infty)\)

OpenStudy (jhannybean):

y-0 = 1(x+3) y = x + 3 It's the same thing, your first equation is in slope-intercept form and this equation is in point-slope form.

OpenStudy (anonymous):

so for all of my equations the domain and range would have no restrictions?

OpenStudy (jhannybean):

Nope.

OpenStudy (anonymous):

oh.. how do i know when there is a restriction?

OpenStudy (jhannybean):

When there is either a removable discontinuity, (your function y is in some kind of fraction form where you have to solve for the denominator and the numerator) In these cases you'll be finding vertical and horizontal asymptotes from my knowledge of it, or from what I can remember.

OpenStudy (anonymous):

Lol... can you rephrase that in baby words please?

OpenStudy (anonymous):

im sorry i really dont understand

OpenStudy (jhannybean):

Ok, like... if your function was \[y=\sqrt{x}\] Your restriction for the domain would be \(x \ge 0\)

OpenStudy (jhannybean):

That is a restriction because it "restricts" your domain to being only real numbers ranging from 0 to +ve infinity.

OpenStudy (anonymous):

so anything in the negitives would be wrong..?

OpenStudy (jhannybean):

It wouldn't be wrong, you would get imaginary values, domain restrictions with i's and stuff.

OpenStudy (anonymous):

okay.. thank you soooo much for you help but i honestly dont think im ever going to understand this. i cant grasp it.

OpenStudy (jhannybean):

Try this. http://www.sparknotes.com/math/algebra2/functions/section2.rhtml